All firms in a competitive industry have the following (firm-level) long-run total cost curve:

C(q) = q3–10q2 + 36q

where q is the output of the firm.
a. Compute the long run equilibrium price. What does the long-run supply curve look like if this is a constant cost industry? Explain.
b. Suppose the market demand is given by Q = 111–p. Determine the long-run equilibrium number of firms in the industry.
Answer #1

C(q) = q² 109² +362 q = -log +36 Lury Run Equilibrium Price = min Ad ACE dal zo dy 29-1020 q=5 AC = 25-so+16=11= Equilibsium